OpenStudy (anonymous):

Which is the simplified form of 16 times q to the fifth power times r cubed all over 2 times q to the fourth power times r to the fifth power.?

OpenStudy (anonymous):

\[\frac{ 16q ^{5}r ^{3} }{ 2q ^{4}r ^{5} }\]

OpenStudy (anonymous):

\[\frac{ 8q }{ r ^{2}} \frac{ 14r ^{2} }{ q} \frac{ 8r ^{2} }{ q } \frac{ 14q }{ r ^{2} }\]

OpenStudy (gabylovesyou):

you dont have 2q/r^2 ??

OpenStudy (anonymous):

nope

OpenStudy (gabylovesyou):

16/2 = 8 r^3 divided by r^5 = r^(-2) = 1/r^2 q^5 divided by q^4 = q^1 = q

OpenStudy (gabylovesyou):

so the answer is 8q/r^2

OpenStudy (anonymous):

thank you so much! can you help me with a few others?

OpenStudy (gabylovesyou):

i can try my best lol